OpenStudy (anonymous):

A thermometer is removed from a room where the temperature is 70 F and is taken outside where the air temperature is 10 F. After one half minute the thermometer reds 50 F. what i the reading of the thermometer at t=1min? how long will it take for the thermometer to reach 15 F

OpenStudy (anonymous):

I just need to be sure the givens are correct in other words, \[T_m=10 F\] \[T(0)=70F\] does time have to be in a certain measurement : hours, minutes, seconds?

OpenStudy (anonymous):

looks like newton's law of cooling? use the same time units the problem uses... minutes.

OpenStudy (anonymous):

so T(.5)=50 T(1)=?

OpenStudy (anonymous):

\[\frac{dT}{dt}=k(T-10)\]

OpenStudy (anonymous):

\[\frac{dT}{T-10}=kdt\]

OpenStudy (anonymous):

ln(T-10)=kt+c \[T-10=ce^{kt}\] \[T=10+ce^{kt}\]

OpenStudy (anonymous):

70=10+c c=60 \[T=10+60e^{kt}\]

OpenStudy (anonymous):

\[50=10+60e^{.5k}\] \[40=60e^{.5k}\] \[\frac{40}{60}=e^{.5k}\] \[ln(\frac{1}{3})=.5k\] \[\frac{-ln(3)}{.5}=k\]

OpenStudy (anonymous):

\[T=10+60e^{-2.2}\]

OpenStudy (anonymous):

2/3's wups

OpenStudy (anonymous):

yay error fixed